OpenStudy (lgbasallote):

find the particular solution for \[xyy' = 1 + y^2\] when x = 2; y = 3 the book says it is \(y = \frac{1}{2} \sqrt{ 10x^2 - 4}\) and idk how it got that

OpenStudy (lgbasallote):

i got \[\large \frac{1}{2} \ln (1 + y^2) = \ln x + \frac{\sqrt{10}}{2}\]

OpenStudy (experimentx):

\[ \int \frac{y}{1 + y^2} dy = \int \frac{dx}{x}\] \[ \ln \sqrt{1+y^2} = \ln x + \ln C\] \[ \sqrt{1+ y^2} = xC\] \[ y = +- \sqrt{Cx^2 - 1} \] How do you write +-

OpenStudy (lgbasallote):

\pm

OpenStudy (experimentx):

find the value of C .... using initial values given!! \[ y = \pm \sqrt{Cx^2 - 1} \]

OpenStudy (lgbasallote):

so it would be \[\large y = \pm \sqrt{\frac{\sqrt{10}}{2} x^2 - 1}\] how did that become 10x^2 - 4?

OpenStudy (lgbasallote):

btw..do i always express it as y equals?

OpenStudy (experimentx):

C = 10/4 <--- you don't need to take root

OpenStudy (lgbasallote):

how did you get 10/4?

OpenStudy (experimentx):

also in your answer you should be getting \[ C = \ln \frac{\sqrt{10}}{2}\] cancelling and isolating y you should be able to get the same

OpenStudy (lgbasallote):

\[\large \sqrt{1+y^2} = Cx\] \[\sqrt{1+ (3)^2} = c(2)\] \[\sqrt{10} = 2C\]

OpenStudy (experimentx):

C <-- integral constants are stupid ... try taking constants such that i helps to simplify your equation or expression so C = sqrt{10} by 2 just simplify that equation in terms of y,you will get the same ...

OpenStudy (lgbasallote):

how do i simplify in terms of y?

OpenStudy (lgbasallote):

uhmm okay i got it lol

OpenStudy (lgbasallote):

C^2 is still C

OpenStudy (lgbasallote):

anyway how do you turn \[\sqrt{\frac{10}{4} x^2 - 1}\] into 10x^2 - 4? no denominators?

OpenStudy (experimentx):

\[ \sqrt{1 + y^2} = \frac{\sqrt{10}}{2}x\] square both sides \[ 1+y^2 = \frac{10x^2}{4}\] \[y^2 = \frac{10x^2 - 4}{4} \] take square root ...

OpenStudy (lgbasallote):

oh nevermind i saw it lol...just one question...do i always put it in terms of y equals?

OpenStudy (experimentx):

not necessarily ... there might be some equation whose form cannot be expressed simply interms of y ... leave those. and besides there are not functions ... just solution curves

OpenStudy (experimentx):

*these <-- there

OpenStudy (lgbasallote):

i see thanks

OpenStudy (experimentx):

yw